OpenStudy (anonymous):

Part a. Write the equation of the line that is parallel to the line y = -2/5x and contains the point (5,-3). Part b. Write the equation of a line that is perpendicular to the line y = 1/4x + 2 and contains the point (11,4).

OpenStudy (anonymous):

If the equation has a parallel line then the slope m=-2/5 is the same for the equation. You have the slope, now just use the equation: y-y1=m(x-x1) Where you replace the values, y-(-3)=-2/5(x-5) And you get, y+3=-2/5x+2 Subtract 3, and you get y=-2/5x-1. That is your a. Now do the same for b, except the slope m is the inverse of m=1/4;

OpenStudy (anonymous):

Thank you @gezimbasha

OpenStudy (anonymous):

You're welcome. Do you need help to solve the other problem?

OpenStudy (anonymous):

No, that was the end of it. thank you very much.(: